# MoEMeta: Mixture-of-Experts Meta Learning for Few-Shot Relational Learning

### v1

@@ Abstract
Few-shot knowledge graph relational learning seeks to perform reasoning over relations given only a limited number of training examples. While existing approaches largely adopt a meta-learning framework for enabling fast adaptation to new relations, they suffer from two key pitfalls. First, they learn relation meta-knowledge in isolation, failing to capture common relational patterns shared across tasks. Second, they struggle to effectively incorporate local, task-specific contexts crucial for rapid adaptation. To address these limitations, we propose MoEMeta, a novel meta-learning framework that disentangles globally shared knowledge from task-specific contexts to enable both effective generalization and rapid adaptation. MoEMeta introduces two key innovations: (i) a mixture-of-experts (MoE) model that learns globally shared relational prototypes to enhance generalization, and (ii) a task-tailored adaptation mechanism that captures local contexts for fast task-specific adaptation. By balancing global generalization with local adaptability, MoEMeta significantly advances few-shot relational learning. Extensive experiments and analyses on three KG benchmarks demonstrate that MoEMeta consistently outperforms existing baselines, achieving state-of-the-art performance. [^1] Code is available at: https://github.com/alexhw15/MoEMeta.git.
@@ 1 Introduction
Knowledge graphs (KGs) encode real-world knowledge in the form of factual triplets $(h,r,t)$, where each triplet represents a relationship $r$ between a head entity $h$ and a tail entity $t$ (Bollacker et al., 2008; Suchanek et al., 2007; Vrandečić and Krötzsch, 2014). KGs are crucial for powering various intelligent applications, such as question answering (Wu et al., 2016), Web search (Fang et al., 2017), and advanced recommendation systems (Cao et al., 2019). Despite their widespread adoption, KGs are inherently incomplete. Although numerous KG completion methods  (Bordes et al., 2013; Yang et al., 2015; Trouillon et al., 2016) have been proposed to infer missing facts based on observed triplets, their effectiveness is severely hindered by data sparsity, especially for long-tail or emerging relations with only a handful of known training triplets.
In response, few-shot relational learning (FSRL) has emerged as a promising direction for reasoning over few-shot relations with limited training triplets, enabling generalization to novel relations in data-scarce scenarios. Most existing FSRL methods  (Chen et al., 2019; Jiang et al., 2021; Niu et al., 2021; Wu et al., 2023) build upon model-agnostic meta-learning (MAML) (Finn et al., 2017), a gradient-based meta-learning framework that learns a set of global parameters serving as an initialization for rapid task adaptation using only one or a few steps of gradient descent updates. In FSRL, each task is relation-centric, comprising only a few training triplets of a single relation, and the objective is to predict instances of novel relations not seen during meta-training. These methods focus on learning and encoding meta-knowledge across tasks to support rapid adaptation with minimal supervision.
@@
Second, current methods are limited by their reliance on a single set of global parameters to encode meta-knowledge. With a shared initialization, a uniform adaptation process struggles to capture the unique characteristics of individual tasks from only a few support triplets. This is critical because relations and entities in KGs exhibit diverse interaction patterns (Wang et al., 2014; Ji et al., 2015), such as one-to-one, one-to-many, many-to-one, and many-to-many. An entity often reveals entirely different aspects depending on the relational context. For instance, the entity Elon Musk participates in the CeoOf relation (a professional role) and the FatherOfPerson relation (a family role). Current methods, constrained by a shared starting point, cannot effectively adapt to these locally distant contexts. Failing to modulate meta-knowledge for a specific task results in suboptimal adaptation.
Present work and contributions. To address these limitations, we propose MoEMeta, a novel meta-learning framework to address a critical, non-trivial challenge unaddressed by prior FSRL work: how to disentangle globally shared knowledge from task-specific contexts to achieve effective generalization and rapid adaptation under few-shot settings. MoEMeta provides a principled solution through two novel components that work in synergy.
@@
Together, these two innovations achieve a good balance between global knowledge generalization and fine-grained task-specific adaptation, which is critical for FSRL with limited supervision. Experimental results demonstrate that MoEMeta achieves state-of-the-art performance on three KG benchmarks, significantly outperforming existing methods. In-depth ablation studies and visualizations further validate the efficacy of each component in MoEMeta.
@@ 2 Related Work
Most mainstream FSRL methods are based on meta-learning (Finn et al., 2017), which learns an initial meta-prior from existing tasks for rapid adaptation to new few-shot relations. MetaR (Chen et al., 2019) learns relation-specific meta-information by simply averaging support triplet representations. MetaP (Jiang et al., 2021) utilizes meta-pattern learning under one-shot settings. GANA (Niu et al., 2021) integrates meta-learning with TransH (Wang et al., 2014) and devises a gated and attentive neighbor aggregator to address sparse neighbors. HiRe (Wu et al., 2023) exploits multi-granularity relational information to learn more generalizable meta-knowledge. RelAdapter (Ran et al., 2024) adopts a feedforward network with a residual layer to adapt relation-meta to target relation. Despite their progress, these methods face two key limitations: (1) they often neglect common relational patterns when learning meta-knowledge during meta-training, and (2) they fail to effectively incorporate local, task-specific contexts essential for rapid adaptation. Our work addresses these critical gaps by proposing a principled meta-learning framework that disentangles globally shared knowledge from task-specific adaptation, enabling both effective generalization and fast adaptation in FSRL. For a more extensive review of related work, including KG embedding learning and general few-shot learning techniques, please refer to Appendix C.
@@
The mixture-of-experts (MoE) architecture, originally introduced by Jacobs et al. (1991), is a modular machine learning framework designed to improve performance by dynamically combining multiple specialized models, known as “experts”. Sparse MoEs and their “soft” variants (Jordan and Jacobs, 1994; Puigcerver et al., 2024) have been widely adopted in deep learning, especially in large-scale models for natural language processing and computer vision, to improve model training efficiency and scalability through adaptive subnetwork selection or dynamic pruning (Shazeer et al., 2017; Du et al., 2022). In a related KG context, MoE has been applied to fuse different semantics of relations (structural ontology and textual descriptions) at the triplet level (Song et al., 2022). However, our work innovatively proposes MoE as a meta-learner for FSRL, operating at the task level to learn specialized experts that can capture diverse relational patterns, thereby achieving superior task adaptation and knowledge transfer.
@@ 3.1 FSRL Problem Formulation
A knowledge graph (KG) can be represented as a collection of factual triplets $\mathcal{G}=\{(h,r,t)\in\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$, where $\mathcal{E}$ denotes the set of entities, and $\mathcal{R}$ denotes the set of relations. Each triplet $(h,r,t)$ indicates a relationship $r$ between head entity $h$ and tail entity $t$.
Given a novel relation $r\notin\mathcal{R}$ and its corresponding support set $\mathcal{S}_{r}=\{(h_{i},r,t_{i})\}_{i=1}^{K}$, where $h_{i},t_{i}\in\mathcal{E}$. The objective of few-shot relational learning (FSRL) is to predict the missing tail entities for a query set $\mathcal{Q}_{r}=\{(h_{j},r,?)\}$, given a head entity $h_{j}\in\mathcal{E}$ and a relation $r$. A candidate set $\mathcal{C}_{h_{j},r}$ is provided for each query triplet $(h_{j},r,?)\in\mathcal{Q}_{r}$. The goal is to rank the true tail entity highest among all candidates. The support set $\mathcal{S}_{r}$ and query set $\mathcal{Q}_{r}$ jointly define a task $\mathcal{T}_{r}=(\mathcal{S}_{r},\mathcal{Q}_{r})$ for relation $r$. In $K$-shot relation learning, the support set $\mathcal{S}_{r}$ contains only a few training triplets ($K=1,3,5$).
@@ 3.2 Meta-Learning based FSRL
Given the strength of meta-learning in enabling rapid adaptation with minimal data, we build our approach on a meta-learning based framework, involving two stages: meta-training and meta-testing. The model is expected to learn a meta-prior from seen tasks in the meta-training stage, which is then effectively adapted to novel tasks in the meta-testing stage.
Meta-Training Stage. During meta-training, an FSRL model is trained on a set of relations $\mathcal{R}_{\text{tr}}$ with corresponding task data $\mathcal{D}_{\text{tr}}=\{\mathcal{T}_{r}\mid r\in\mathcal{R}_{\text{tr}}\}$. The model aims to learn a prior, which initializes a relation-meta learner for a new task. For each task $\mathcal{T}_{r}=(\mathcal{S}_{r},\mathcal{Q}_{r})\in\mathcal{D}_{\text{tr}}$, the learner extracts a task-specific relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$ based on the support set $\mathcal{S}_{r}$, which is refined using a gradient meta $\mathbf{G}_{\mathcal{T}_{r}}$ computed from the support loss:
@@
where $\alpha$ is the learning rate. The adapted relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$ is then applied to the query set with the ground-truth tail entities, where the query loss is backpropagated to update both the prior and the entity embedding matrix.
Meta-Testing Stage. During meta-testing, the model is tested on novel relations $\mathcal{R}_{\text{te}}$ and their associated task data $\mathcal{D}_{\text{te}}=\{\mathcal{T}_{r}\mid r\in\mathcal{R}_{\text{te}}\}$, where $\mathcal{R}_{\text{tr}}\cap\mathcal{R}_{\text{te}}=\emptyset$. For each novel task $\mathcal{T}_{r}=(\mathcal{S}_{r},\mathcal{Q}_{r})\in\mathcal{D}_{\text{te}}$, the model utilizes the learned prior to initialize the relation-meta learner and compute an initial relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$, which is then adapted based on the support set $\mathcal{S}_{r}$ to obtain $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$. Finally, for each query triplet $(h,r,?)$, the model ranks all candidate tail entities $t$ in the candidate set $\mathcal{C}_{h,r}$ by computing a score $\textrm{s}(h,\mathbf{R}_{\mathcal{T}_{r}},t)$ for each candidate and sorting them in ascending order.
@@ 4 The Proposed MoEMeta Framework
This section presents our proposed MoEMeta framework. Its design is underpinned by two key insights: First, relations in real-world KGs exhibit heterogeneous characteristics, with some subsets of relations sharing greater similarity and stronger interconnections; Second, relations in KGs manifest diverse interaction patterns, such as one-to-many, many-to-one, and many-to-many. Building upon these insights, we introduce MoEMeta, a novel meta-learning framework that (1) utilizes an MoE model to dynamically learn relational prototypes and generate relation-meta for each task; (2) incorporates a task-tailored adaptation mechanism to facilitate local adaptation to each task.
@@ 4.1 Attentive Neighbor Aggregation
For a given entity $e$, its one-hop neighbors are represented as a set of neighboring tuples $(r_{i},e^{\prime}_{i})\}_{i=1}^{n}$, where $n$ is the total number of neighbors, $e^{\prime}_{i}$ is the neighboring entity and $r_{i}$ denotes the relation between $e^{\prime}_{i}$ and $e$. We first concatenate the embeddings of $r_{i}$ and $e^{\prime}_{i}$ to form the embedding $\mathbf{c}_{i}$ of the $i$-th tuple, i.e.  $\mathbf{c}_{i}=\mathbf{r}_{i}\oplus\mathbf{e}^{\prime}_{i}$, where $\oplus$ denotes concatenation of two vectors. To capture interactions among neighbors, we then apply a linear transformation followed by a non-linear activation function:
@@
where $\mathbf{W}$ is a learnable weight matrix. To enable the model to attend to informative neighbors, we compute an attention-like gate value $g_{i}$ for each neighboring tuple using a sigmoid function:
@@
$\mathbf{e}=\text{Aggregate}(\{g_{i}\cdot\mathbf{c}^{\prime}_{i}\}_{i=1}^{n})+\mathbf{e},$ (4)
The neighbor encoder is applied to both head entity $h$ and tail entity $t$ independently to ensure that the final embedding reflects both the entity’s intrinsic information and its local neighborhood structure.
@@ 4.2 MoE-based Meta-Knowledge Learning
To overcome the limitation of existing methods that overlook shared meta-knowledge across related tasks, we introduce a globally learned MoE model that captures common relational patterns, termed relational prototypes. For each task, a subset of relevant experts is dynamically selected from a shared pool based on its support triplets to learn relation-meta. This mechanism enables adaptive relation decomposition and expert specialization, facilitating the learning of diverse, compositional relational patterns and promoting effective generalization across tasks.
Specifically, we design a sparsely-gated MoE model consisting of $M$ experts. Each expert within the MoE is architecturally identical to a standard MLP (multilayer perceptron). For a given support triplet, a gating mechanism dynamically routes the triplet to a selected subset of experts, assigning weights to each expert in the subset based on its relevance. The outputs of the top-$N$ experts are then aggregated according to their weights, resulting in a representation of the relation associated with the triplet. To compute the relation-meta for a support set, we average the relation representations derived from all triplets within the support set. This approach enables the MoE layer to retrieve the most relevant meta-knowledge from the shared expert pool to model diverse relational patterns.
@@
where $M$ denotes the total number of experts, each implemented as a two-layer MLP, $\text{MLP}_{i}(\cdot;\bm{\theta}_{i})$. For each support triplet $j$, a gate network, $\text{Gate}(\cdot;\bm{\theta}_{g})$, which is also a two-layer MLP, computes a relevance score $s_{i,j}$ for each expert $i$. These scores are normalized via a softmax function over all $M$ experts. The function $\text{Top}_{j}(\cdot,N)$ selects only the top $N$ highest scores among all $M$ experts for each triplet $j$ and setting the rest to zero. This sparsity mechanism enables the model to selectively activate only a subset of experts for each support triplet, thereby facilitating the learning of more specialized relational prototypes globally shared across tasks.
@@
$\mathbf{R}_{\mathcal{T}_{r}}=\frac{1}{K}\textstyle\sum_{i=1}^{K}\mathbf{r}_{i}$ (8)

### v2

@@ Abstract
Few-shot knowledge graph relational learning seeks to perform reasoning over relations given only a limited number of training examples. While existing approaches largely adopt a meta-learning framework for enabling fast adaptation to new relations, they suffer from two key pitfalls. First, they learn relation meta-knowledge in isolation, failing to capture common relational patterns shared across tasks. Second, they struggle to effectively incorporate local, task-specific contexts crucial for rapid adaptation. To address these limitations, we propose MoEMeta, a novel meta-learning framework that disentangles globally shared knowledge from task-specific contexts to enable both effective model generalization and rapid adaptation. MoEMeta introduces two key innovations: (i) a mixture-of-experts (MoE) model that learns globally shared relational prototypes to enhance generalization, and (ii) a task-tailored adaptation mechanism that captures local contexts for fast task-specific adaptation. By balancing global generalization with local adaptability, MoEMeta significantly advances few-shot relational learning. Extensive experiments and analyses on three KG benchmarks show that MoEMeta consistently outperforms existing baselines, achieving state-of-the-art performance. [^1] The code is available at: https://github.com/alexhw15/MoEMeta.
@@ 1 Introduction
Knowledge graphs (KGs) encode real-world knowledge in the form of factual triplets $(h,r,t)$, where each triplet represents a relationship $r$ between a head entity $h$ and a tail entity $t$ (Bollacker et al., 2008; Suchanek et al., 2007; Vrandečić and Krötzsch, 2014).KGs are crucial for powering various intelligent applications, such as question answering (Wu et al., 2016), Web search (Fang et al., 2017), and advanced recommendation systems (Cao et al., 2019). Despite their widespread adoption, real-world KGs are inherently incomplete. Although numerous KG completion methods  (Bordes et al., 2013; Yang et al., 2015; Trouillon et al., 2016) have been proposed to infer missing facts based on observed triplets, their effectiveness is severely hindered by data sparsity, especially for long-tail or emerging relations with only a handful of known training triplets.
In response, few-shot relational learning (FSRL) has emerged as a promising direction for reasoning over few-shot relations with limited training triplets, enabling generalization to novel relations in data-scarce scenarios. Most existing FSRL methods  (Chen et al., 2019; Jiang et al., 2021; Niu et al., 2021; Wu et al., 2023) are built upon model-agnostic meta-learning (MAML) (Finn et al., 2017), a gradient-based meta-learning framework that learns a set of global parameters as an initialization for rapid task adaptation using only one or a few steps of gradient descent updates. In FSRL, each task is relation-centric, comprising only a few training triplets of a single relation, and the objective is to predict instances of novel relations not seen during meta-training. These methods focus on learning and encoding meta-knowledge among tasks to support rapid adaptation with minimal supervision.
@@
Second, current methods are limited by their reliance on a single set of global parameters to encode meta-knowledge. With a shared initialization, a uniform adaptation process lacks the flexibility to capture the unique characteristics of individual tasks from only a few support triplets. This limitation is critical because relations and entities in KGs exhibit diverse interaction patterns (Wang et al., 2014; Ji et al., 2015), such as one-to-one, one-to-many, many-to-one, and many-to-many. An entity often reveals entirely different aspects depending on its relational context. For example, the entity Elon Musk participates in the CeoOf relation (a professional role) and the FatherOfPerson relation (a family role). Methods constrained by a shared initialization struggle to adapt to such locally distant contexts, and failing to modulate meta-knowledge for a specific task results in suboptimal adaptation.
Present work and contributions. To overcome these limitations, we propose MoEMeta, a novel meta-learning framework to tackle a critical, non-trivial challenge not addressed by prior FSRL work: how to disentangle globally shared knowledge from task-specific contexts to achieve effective generalization and rapid adaptation under few-shot settings. MoEMeta provides a principled solution through two novel components that work in synergy.
@@
Together, these two innovations strike an effective balance between global knowledge generalization and fine-grained task-specific adaptation, which is crucial for FSRL with limited supervision. Experimental results demonstrate that MoEMeta achieves state-of-the-art performance on three KG benchmarks, significantly outperforming existing methods. In-depth ablation studies and visualizations further validate the efficacy of each component in MoEMeta.
@@ 2 Related Work
Most mainstream FSRL methods are based on meta-learning (Finn et al., 2017), which learns an initial meta-prior from existing tasks for rapid adaptation to new few-shot relations. MetaR (Chen et al., 2019) learns relation-specific meta-information by simply averaging support triplet representations. MetaP (Jiang et al., 2021) utilizes meta-pattern learning under one-shot settings. GANA (Niu et al., 2021) integrates meta-learning with TransH (Wang et al., 2014) and devises a gated, attentive neighbor aggregator to handle sparse neighborhood information. HiRe (Wu et al., 2023) exploits multi-granularity relational information to learn more generalizable meta-knowledge. RelAdapter (Ran et al., 2024) adopts a shallow feedforward network with a residual connection to adapt relation meta-knowledge within each target task. Despite their progress, these methods face two key limitations: (1) they often neglect common relational patterns when learning meta-knowledge during meta-training, and (2) they fail to effectively incorporate local, task-specific contexts essential for rapid adaptation. Our work addresses these critical gaps by proposing a principled meta-learning framework that disentangles globally shared knowledge from task-specific adaptation, enabling both effective generalization and fast adaptation in FSRL. For a more extensive review of related work, including KG embedding learning and general few-shot learning techniques, please refer to Appendix LABEL:app:related-works.
@@
The mixture-of-experts (MoE) architecture, originally introduced by Jacobs et al. (1991), is a modular machine learning framework designed to improve performance by dynamically combining multiple specialized models, known as “experts”. Sparse MoEs and their “soft” variants (Jordan and Jacobs, 1994; Puigcerver et al., 2024) have been widely adopted in deep learning, especially in large-scale models for natural language processing and computer vision, to improve model training efficiency and scalability through adaptive subnetwork selection or dynamic pruning (Shazeer et al., 2017; Du et al., 2022). In a related KG context, MoE has been applied to fuse different semantics of relations (structural ontology and textual descriptions) at the triplet level (Song et al., 2022). In contrast, our work innovatively proposes MoE as a meta-learner for FSRL that operates at the task level to learn specialized experts capturing diverse relational patterns, thereby enabling superior cross-task knowledge transfer and task adaptation.
@@ 3.1 FSRL Problem Formulation
A knowledge graph (KG) can be represented as a collection of factual triplets $\mathcal{G}=\{(h,r,t)\in\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$, where $\mathcal{E}$ denotes the set of entities, and $\mathcal{R}$ denotes the set of relations. Each triplet $(h,r,t)$ indicates a relationship $r$ between the head entity $h$ and the tail entity $t$.
Given a novel relation $r\notin\mathcal{R}$ and its corresponding support set $\mathcal{S}_{r}=\{(h_{i},r,t_{i})\}_{i=1}^{K}$, where $h_{i},t_{i}\in\mathcal{E}$, the objective of few-shot relational learning (FSRL) is to predict the missing tail entities in a query set $\mathcal{Q}_{r}=\{(h_{j},r,?)\}$, given a head entity $h_{j}\in\mathcal{E}$ and a relation $r$. For each query triplet $(h_{j},r,?)\in\mathcal{Q}_{r}$, a candidate set $\mathcal{C}_{h_{j},r}$ is provided, and the goal is to rank the true tail entity highest among all candidates. The support set $\mathcal{S}_{r}$ and the query set $\mathcal{Q}_{r}$ jointly define a task $\mathcal{T}_{r}=(\mathcal{S}_{r},\mathcal{Q}_{r})$ for a relation $r$. In $K$-shot relation learning, the support set $\mathcal{S}_{r}$ contains only a few training triplets (typically $K=1,3,5$).
@@ 3.2 Meta-Learning based FSRL
Given the strength of meta-learning in enabling rapid adaptation with minimal data, we build our approach upon a meta-learning based framework, involving two stages: meta-training and meta-testing. The model is expected to learn a meta-prior from seen tasks in the meta-training stage, which can then be adapted to novel tasks during the meta-testing stage.
Meta-Training Stage. During meta-training, an FSRL model is trained on a set of relations $\mathcal{R}_{\text{tr}}$ with the corresponding task data $\mathcal{D}_{\text{tr}}=\{\mathcal{T}_{r}\mid r\in\mathcal{R}_{\text{tr}}\}$. The model aims to learn a prior, which initializes a relation-meta learner for a new task. For each task $\mathcal{T}_{r}=(\mathcal{S}_{r},\mathcal{Q}_{r})\in\mathcal{D}_{\text{tr}}$, the learner extracts a task-specific relation-meta $\mathbf{R}_{\mathcal{T}_{r}}\in\mathbb{R}^{d}$ based on the support set $\mathcal{S}_{r}$, which is rapidly updated using a gradient meta $\mathbf{G}_{\mathcal{T}_{r}}$ computed from the loss on the support set:
@@
where $\alpha$ is the learning rate. The updated relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$ is then applied to the query set with the ground-truth tail entities, where the query loss is backpropagated to update both the prior and the entity embedding matrix.
Meta-Testing Stage. During meta-testing, the model is evaluated on novel relations $\mathcal{R}_{\text{te}}$ and their associated task data $\mathcal{D}_{\text{te}}=\{\mathcal{T}_{r}\mid r\in\mathcal{R}_{\text{te}}\}$, where $\mathcal{R}_{\text{tr}}\cap\mathcal{R}_{\text{te}}=\emptyset$. For each novel task $\mathcal{T}_{r}=(\mathcal{S}_{r},\mathcal{Q}_{r})\in\mathcal{D}_{\text{te}}$, the model uses the learned prior to initialize the relation-meta learner and compute an initial relation-meta $\mathbf{R}_{\mathcal{T}_{r}}$, which is then adapted using its support set $\mathcal{S}_{r}$ to obtain $\mathbf{R}_{\mathcal{T}_{r}}^{\prime}$. Finally, for each query triplet $(h,r,?)$, the model ranks all candidate tail entities $t$ in the candidate set $\mathcal{C}_{h,r}$ by computing a score $\textrm{s}(h,\mathbf{R}_{\mathcal{T}_{r}}^{\prime},t)$ for each candidate $t\in\mathcal{C}_{h,r}$ and sorting them in ascending order.
@@ 4 The Proposed MoEMeta Framework
This section presents our proposed MoEMeta framework. Its design is underpinned by two key insights: First, relations in real-world KGs exhibit heterogeneous characteristics, with some subsets of relations sharing greater similarity and stronger interconnections; Second, relations in KGs manifest diverse interaction patterns, such as one-to-many, many-to-one, and many-to-many. Building upon these insights, we introduce MoEMeta, a novel meta-learning framework that (1) utilizes an MoE model to dynamically learn relational prototypes and generate relation-meta for each task; and (2) incorporates a task-tailored adaptation mechanism to facilitate local adaptation to each task.
@@ 4.1 Attentive Neighbor Aggregation
For a given entity $e$, its one-hop neighbors are represented as a set of neighboring tuples $(r_{i},e^{\prime}_{i})\}_{i=1}^{n}$, where $n$ is the total number of neighbors, $e^{\prime}_{i}$ is the neighboring entity and $r_{i}$ denotes the relation between $e^{\prime}_{i}$ and $e$. We first concatenate the embeddings of $r_{i}$ and $e^{\prime}_{i}$ to form the embedding of the $i$-th tuple, $\mathbf{c}_{i}=\mathbf{r}_{i}\oplus\mathbf{e}^{\prime}_{i}$, where $\oplus$ denotes concatenation of two vectors. To capture interactions among neighbors, we then apply a linear transformation followed by a non-linear activation function:
@@
where $\mathbf{W}$ is a learnable weight matrix. To enable the encoder to attend to informative neighbors, we compute an attention-like gate value $g_{i}$ for each neighboring tuple using a sigmoid function:
@@
$\mathbf{e}=\text{Aggregate}(\{g_{i}\cdot\mathbf{c}^{\prime}_{i}\}_{i=1}^{n})+\mathbf{e}.$ (4)
The neighbor encoder is applied to head entity $h$ and tail entity $t$ independently to ensure that the final embedding reflects both the entity’s intrinsic information and its local neighborhood structure.
@@ 4.2 MoE-based Meta-Knowledge Learning
To overcome the limitation of existing methods that neglect shared meta-knowledge across related tasks, we introduce a globally learned MoE model that captures common relational patterns, termed relational prototypes. For each task, a subset of relevant experts is dynamically selected from a shared pool based on its support triplets to learn the relation-meta. This mechanism enables adaptive relation decomposition and expert specialization, facilitating the learning of diverse, compositional relational patterns and promoting effective generalization across tasks.
Specifically, we design a sparsely-gated MoE model consisting of $M$ experts. Each expert within the MoE is architecturally identical to a standard MLP (multilayer perceptron). For a given support triplet, a gating mechanism dynamically routes the triplet to a selected subset of experts, assigning weights to each expert in the subset based on its relevance. The outputs of the top-$N$ experts are then aggregated according to their weights, resulting in a representation of the relation associated with the triplet. To compute the relation-meta for a support set, we average the relation representations derived from all triplets within the support set. This approach enables the MoE to retrieve the most relevant meta-knowledge from the shared expert pool to model diverse relational patterns.
@@
where $M$ denotes the total number of experts, each implemented as a two-layer MLP, $\text{MLP}_{i}(\cdot;\bm{\theta}_{i})$. For each support triplet $j$, a gate network, $\text{Gate}(\cdot;\bm{\theta}_{g})$, also implemented as a two-layer MLP, computes a relevance score $s_{i,j}$ for each expert $i$. These scores are normalized via a softmax function over all $M$ experts. The function $\text{Top}_{j}(\cdot,N)$ selects only the top $N$ highest scores among all $M$ experts for each triplet $j$, setting the remaining scores to zero. This sparsity mechanism enables the model to selectively activate only a subset of experts for each support triplet, thereby facilitating the learning of more specialized relational prototypes globally shared across tasks.
@@
$\mathbf{R}_{\mathcal{T}_{r}}=\frac{1}{K}\sum_{i=1}^{K}\mathbf{r}_{i}$ (8)